# Generalization bounds for mixing processes via delayed online-to-PAC conversions

### v1

@@ 3 Proving generalization bounds via online learning
Theorem 3 (Bound in probability).
Assume that $(Z_{t})_{t\in\mathbb{N}^{*}}$ satisfies Assumption 1 and consider a $d$-delayed online learning algorithm with regret bounded by $R_{d,n}(P^{*})$ against any comparator $P^{*}$. Then, for any $\delta>0$, it holds with probability $1-\delta$ on the draw of $S_{n}$, uniformly for all $\mathcal{A}$,
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{R_{d,n}(P_{W_{n}|S_{n}})}{n}+\phi_{d}+\sqrt{\frac{2d\log\left(\frac{d}{\delta}\right)}{n}}.$
The proof of this claim follows directly from combining the decomposition of Theorem 1 with a standard concentration result for mixing processes that we state below.
Lemma 2.
Fix $d\in[\![1,n]\!]$ and consider $(Z_{t})_{t\in\mathbb{N}^{*}}$ satisfying Assumption 1. Consider the generalization game of Definition 1. Then, for any $\delta>0$, the following bound is satisfied with probability at least $1-\delta$:
$M_{n}\leq\phi_{d}+\sqrt{\frac{2d\log\left(\frac{d}{\delta}\right)}{n}}.$
The proof is based on a classic “blocking” technique due to Yu, (1994). For the sake of completeness, we provide a proof in Appendix A.2.
@@
The dependence on the delay $d$ for the bounds that we presented in the previous section is non-trivial. Indeed, if on the one hand increasing the delay will reduce the magnitude of $\phi_{d}$, on the other hand the regret of the online learner will grow with $d$. There is hence a trade-off between these two terms appearing in our bounds. In what follows, we derive some concrete generalization bounds from Theorem 3, under a number of different choices of the online learning algorithm. For concreteness, we will consider two types of mixing assumptions, but stress that the approach can be applied to any process that satisfies Assumption 1.
@@ 4.1 Regret bounds for delayed online learning
From Theorem 3, we can obtain a generalization bound using our framework if we have a regret bound for a delayed online algorithm. This is a well-known problem in the area of online learning (see, e.g., Weinberger and Ordentlich,, 2002; Joulani et al.,, 2013). In the following, we will leverage the following simple trick that allows us to extend the regret bounds of any online learning algorithm to its delayed counterpart, provided that the regret bound respects some specific assumptions.
Lemma 3 (Weinberger and Ordentlich,, 2002).
Consider any online algorithm whose regret satisfies $\mathrm{Regret}_{n}(P^{*})\leq R(n;P^{*})$ for any comparator $P^{*}$, where $R$ is a non-decreasing real-valued function of its first argument such that $y\mapsto yR(x/y;P^{*})$ is a concave function of $y$ for any fixed $x$ and any $P^{*}$. Then, for any $d\geq 1$ there exists an online learning algorithm with delay $d$ such that, for any comparator $P^{*}$,
$\mathrm{Regret}_{d,n}(P^{*})\leq dR\left(n/d;P^{*}\right)\,.$
The proof idea is closely related to the blocking trick of Yu, (1994), with an algorithmic construction that runs one instance of the base method for each index $i=1,2,\dots,d$, with the $i$-th instance being responsible for the regret in rounds $i,i+d,i+2d,\dots$ (more details are provided in Appendix B.3). For most of the regret bounds that we consider, the function $R$ takes the form $R(n;P^{*})=O(\sqrt{n})$, so that the first term in the generalization bound is typically of order $\sqrt{d/n}$. Since this term matches the bound on $M_{n}$ in Lemma 2, in this case the final generalization bound behaves effectively as if the sample size was $n/d$ instead of $n$.
@@ 4.2 Geometric and algebraic mixing
The following definition gives two concrete examples of mixing processes that satisfy Assumption 1 with different choices of $\phi_{d}$, and are commonly considered in the related literature (see, e.g., Mohri and Rostamizadeh,, 2010, Levin and Peres,, 2017).
Definition 2.
We say that a stationary process $(Z_{t})_{t\in\mathbb{N}^{*}}$ satisfying Assumption 1 is:
  • geometrically mixing, if $\phi_{d}=Ce^{-\frac{d}{\tau}}$, for some positive $\tau$ and $C$;
  • algebraically mixing, if $\phi_{d}=Cd^{-r}$, for some positive $r$ and $C$.
Instantiating the bound of Theorem 3 to these two cases yields the following two corollaries.
Corollary 1.
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})}{n}+\frac{C}{n}+\sqrt{\frac{2\left(\tau\log n+1\right)\log\left(\frac{\tau\log n+1}{\delta}\right)}{n}}\,.$
Up to a term linear in $\tau$ and some logarithmic factors, the above states that under the geometric mixing condition, the same rates are achievable as in the i.i.d. setting. Roughly speaking, this amounts to saying that the effective sample size is a factor $\tau$ smaller than the original number of samples $n$, as long as generalization is concerned.
Corollary 2.
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})}{n}+C\left(1+\sqrt{\log(d/\delta)}\right)n^{-\frac{2r}{2(1+2r)}}\,.$
@@ 4.3 Multiplicative weights with delay
We now turn our attention to picking online strategies for the purpose of bounding the main term in the decomposition of the generalization error. We start by focusing on the classic exponential weighted average (EWA) algorithm (Vovk,, 1990; Littlestone and Warmuth,, 1994; Freund and Schapire,, 1997). We fix a data-free prior $P_{1}\in\Delta_{\mathcal{W}}$ and a learning rate parameter $\eta>0$. We consider the updates
$P_{t+1}=\operatorname*{arg\ min}_{P\in\Delta_{\mathcal{W}}}\left\{\langle P,c_{t}\rangle+\frac{1}{\eta}\mathcal{D}_{\mathrm{KL}}\left(P\middle\|P_{t}\right)\right\},$
Combining the standard regret bound of EWA (see Appendix B.1) with Lemma 3 and Corollary 1 yields the result that follows.
Corollary 3.
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{(\tau\log n+1)\mathcal{D}_{\mathrm{KL}}\left(P^{*}\middle\|P_{1}\right)}{\eta n}+\frac{\eta}{2}+\frac{C}{n}+\sqrt{\frac{2\left(\tau\log n+1\right)\log\left(\frac{\tau\log n+1}{\delta}\right)}{n}}\,.$
This results suggests that when considering geometric mixing processes, by applying a union bound over a well-chosen range of $\eta$ we recover the PAC-Bayes bound of McAllester, (1998) up to the $O(\sqrt{\tau\log n})$ factor that quantifies the price of dropping the i.i.d. assumption. A similar result can be derived from Corollary 2 for algebraically mixing processes, leading to a bound typically scaling as $n^{-2r/(2(1+2r))}$.
@@ 4.4 Follow the regularized leader with delay
Corollary 4.
Suppose that $(Z_{t})_{t\in\mathbb{N}^{*}}$ is a geometric mixing process with constants $\tau,C>0$. Suppose that $\ell(w,z)\in[0,1]$ for all $w,z$. Assume there exists $B>0$ such that for all $t,\left\|c_{t}\right\|_{*}\leq B.$ Then, for any $P_{1}\in\Delta_{\mathcal{W}}$, for any $\delta>0$ with probability at least $1-\delta$ on the draw of $S_{n}$, uniformly for all $\mathcal{A}$,
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\left(h(P^{*})-h(P_{1})\right)(\tau\log n+1)}{\eta n}+\frac{\eta B^{2}}{2\alpha}+\frac{C}{n}+\sqrt{\frac{2\left(\tau\log n+1\right)\log\left(\frac{\tau\log n+1}{\delta}\right)}{n}}.$
This generalization bound is similar to the bound of Theorem 9 of Lugosi and Neu, (2023) up to the $O(\sqrt{\tau\log n})$ factor capturing the effect of the non-i.i.d. data. In particular, this result recovers PAC-Bayesian bounds like those of Corollary 3 when choosing $h=\mathcal{D}_{\mathrm{KL}}\left(\cdot\middle\|P_{1}\right)$. We refer to Section 3.2 in Lugosi and Neu, (2023) for more discussion on such bounds. As before, a similar result can be stated for algebraically mixing processes, with the leading terms approaching zero at rate of $n^{-2r/2(1+2r)}$ instead of $n^{-1/2}$.
@@ 5 Generalization bounds for dynamic hypotheses
Inspired by the works of Eringis et al., (2022, 2024), we also extend our framework to accommodate loss functions $\ell$ that rely not only on the last data point $Z_{t}$, but on the entire data sequence $\overline{Z}_{t}=\left(Z_{t},Z_{t-1},\dots,Z_{1}\right)$. Formally, we will consider loss functions of the form $\ell:\mathcal{W}\times\mathcal{Z}^{*}\rightarrow{}_{+}$ [^1] Here, $\mathcal{Z}^{*}$ denotes the disjoint union $\mathcal{Z}^{*}=\sqcup_{t\in\mathbb{N}}\mathcal{Z}^{t}$. and write $\ell(w,\overline{z}_{t})$ to denote the loss associated with hypothesis $w\in\mathcal{W}$ on sequence $\overline{z}_{t}\in\mathcal{Z}^{t}$. This consideration extends the learning problem to class of dynamical predictors such as Kalman filters, autoregressive models, or recurrent neural networks (RNNs), broadly used in time-series forecasting (Ariyo et al.,, 2014; Takeda et al.,, 2016). Specifically, if we think of $z_{t}=(x_{t},y_{t})$ as a data-pair of context and observation, in time-series prediction we usually not only rely on the context $x_{t}$ but also on the past sequence of contexts and observations $(x_{t-1},y_{t-1},\dots,x_{1},y_{1})$. As an example, consider $\ell(w,z_{t},\dots,z_{1})=\frac{1}{2}(y_{t}-h_{w}(x_{t},z_{t-1},\dots,z_{1}))^{2}$ where $h\in\mathcal{H}$ is a function class parameterized by $\mathcal{W}$. For this type of loss function a natural definition of the test error is:
@@
This is a generalization of Assumption 1 in the sense that taking $\ell(w,Z_{t},\dots,Z_{1})=\ell(w,Z_{t})$ simply amounts to requiring the same mixing condition as before. For our online-to-PAC conversion we consider the same framework as in Definition 1, except that now the cost function is defined as
@@
Then it easy to check that result of Lemma 2 still holds for this specific cost, and we can thus extend all the results of Section 4. For concreteness, we state the following adaptation of Theorem 3 below.
Theorem 4.
Assume $(Z_{t})_{t\in\mathbb{Z}}$ which satisfies Assumption 2 and consider a $d$-delayed online learning algorithm with regret bounded by $R_{d,n}(P^{*})$ against any comparator $P^{*}$. Then, for any $\delta>0$, it holds with probability $1-\delta$:
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{R_{d,n}(P_{W_{n}|S_{n}})}{n}+\phi_{d}+\sqrt{\frac{2d\log\left(\frac{d}{\delta}\right)}{n}}.$
@@
Assumption 3.
There exists a decreasing sequence $(B_{d})_{d\in\mathbb{N}^{*}}$ of non-negative real numbers such that for any two sequences $\overline{z}_{t}=(z_{t},\dots,z_{i})$ and $\overline{z}_{t}^{\prime}=(z_{t}^{\prime},\dots,z_{j}^{\prime})$ of possibly different lengths that satisfy $z_{k}=z_{k}^{\prime}$ for all $k\in{t,\dots,t-d+1}$, we have $\left|\ell(w,\overline{z}_{t})-\ell(w,\overline{z}^{\prime}_{t})\right|\leq B_{d},$ for all $w\in\mathcal{W}$.
This condition can be verified for stable dynamical systems like autoregressive models, certain classes of stable RNNs, or sequential predictors that have bounded memory by design (see Eringis et al.,, 2022, 2024). The next assumption is a refinement of Assumption 1, adapted to the case where the loss function acts on blocks of $d$ data points $\overline{z}_{t-d+1:t}=(z_{t},z_{t-1},\dots,z_{t-d+1})$.
Assumption 4.
Let $\overline{Z}_{t}=(Z_{t},\dots,Z_{1})$ be a sequence of data points and let $\overline{Z}_{t}^{\prime}=(Z_{t}^{\prime},\dots,Z_{0}^{\prime},\dots)$ be an independent copy of the same process. Then, there exists a decreasing sequence $(\beta_{d})_{d\in\mathbb{N}^{*}}$ non-negative real numbers such that the following is satisfied for all hypotheses $w\in\mathcal{W}$ and all $d\in\mathbb{N}^{*}$:
@@
This assumption can be verified whenever the loss function is bounded and the joint distribution of the data block $\overline{Z}_{t-d+1:t}$ satisfies a $\beta$-mixing assumption. In more detail, this latter condition amounts to requiring that the conditional distribution of each data block given a block that trails $d$ steps behind is close to the marginal distribution in total variation distance, up to an additive term of $\beta_{d}$. The following proposition shows that these two simple conditions together imply that Assumption 2 holds, and that thus the bound of Theorem 4 can be meaningfully instantiated for bounded-memory hypothesis classes deployed on mixing processes.
Proposition 2.

### v3

@@ 3 Proving generalization bounds via online learning
Theorem 3.9 (Bound in probability).
Assume that $(Z_{t})_{t\in\mathbb{N}^{*}}$ satisfies Assumption 1 and consider a $d$-delayed online learning algorithm with regret bounded by $\mathrm{Regret}_{d,n}(P^{*})$ against any comparator $P^{*}$. Then, for any $\delta>0$, it holds with probability $1-\delta$ on the draw of $S_{n}$, uniformly for all $\mathcal{A}$,
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})}{n}+\phi_{d}+\sqrt{\frac{2d\log\left(\frac{1}{\delta}\right)}{n}}.$
The proof of this claim follows directly from combining the decomposition of Theorem 3.1 with a standard concentration result for mixing processes that we state below.
Lemma 3.10.
Fix $d\in[\![1,n]\!]$ and consider $(Z_{t})_{t\in\mathbb{N}^{*}}$ satisfying Assumption 1. Consider the generalization game of Definition 3.4. Then, for any $\delta>0$, the following bound is satisfied with probability at least $1-\delta$:
$M_{n}\leq\phi_{d}+\sqrt{\frac{2d\log\left(\frac{1}{\delta}\right)}{n}}.$
The proof is based on a classic “blocking” technique due to Yu (1994). For the sake of completeness, we provide a proof in Appendix A.2.
@@
The dependence on the delay $d$ for the bounds that we presented in the previous section is non-trivial. Indeed, if on the one hand increasing the delay will reduce the magnitude of $\phi_{d}$, on the other hand the regret of the online learner will grow with $d$. There is hence a trade-off between these two terms appearing in our bounds. In what follows, we derive some concrete generalization bounds from Theorem 3.9, under a number of different choices of the online learning algorithm. For concreteness, we will consider two types of mixing assumptions, but stress that the approach can be applied to any process that satisfies Assumption 1.
@@ 4.1 Regret bounds for delayed online learning
From Theorem 3.9, we can obtain a generalization bound using our framework if we have a regret bound for a delayed online algorithm. This is a well-known problem in the area of online learning (see, e.g., Weinberger and Ordentlich, 2002; Joulani et al., 2013). In the following, we will leverage the following simple trick that allows us to extend the regret bounds of any online learning algorithm to its delayed counterpart, provided that the regret bound respects some specific assumptions.
Lemma 4.1 (Weinberger and Ordentlich, 2002).
Consider any online algorithm whose regret satisfies $\mathrm{Regret}_{n}(P^{*})\leq R(n)$ for any comparator $P^{*}$, where $R$ is a non-decreasing real-valued function such that $y\mapsto yR(x/y)$ is concave in $y$ for any fixed $x$. Then, for any $d\geq 1$ there exists an online learning algorithm with delay $d$ such that, for any $P^{*}$,
$\mathrm{Regret}_{d,n}(P^{*})\leq dR\left(n/d\right)\,.$
The proof idea is closely related to the blocking trick of Yu (1994), with an algorithmic construction that runs one instance of the base method for each index $i=1,2,\dots,d$, with the $i$-th instance being responsible for the regret in rounds $i,i+d,i+2d,\dots$ (more details are provided in Appendix B.3). For most of the regret bounds that we consider, the function $R$ takes the form $R(n)=O(\sqrt{n})$, so that the first term in the generalization bound is typically of order $\sqrt{d/n}$. Since this term matches the bound on $M_{n}$ in Lemma 3.10, in this case the final generalization bound behaves effectively as if the sample size was $n/d$ instead of $n$.
We remark that that the regret guarantees of Weinberger and Ordentlich (2002) are minimax optimal (Joulani et al., 2013) [^1] The guarantees of Lemma 4.1 are minimax optimal in the sense that the best that one can hope for without further assumptions on the loss is a regret of $O(\sqrt{dT})$. Although it is true that better regret bounds can be achieved under more specific assumptions about the losses (see Table 1 in Joulani et al., 2013), this is precisely the type of assumption that we wanted to avoid in our work. , and although more efficient methods exist for online learning with delays, the online-to-PAC reduction does not require the strategy to be executed in practice, making the computational efficiency of the online method irrelevant to our analysis.
@@ 4.2 Geometric and algebraic mixing
The following definition gives two concrete examples of mixing processes that satisfy Assumption 1 with different choices of $\phi_{d}$, and are commonly considered in the related literature (see, e.g., Mohri and Rostamizadeh, 2010, Levin and Peres, 2017).
Definition 4.2.
We say that a stationary process $(Z_{t})_{t\in\mathbb{N}^{*}}$ satisfying Assumption 1 is geometrically mixing if $\phi_{d}=Ce^{-\frac{d}{\tau}}$, for some positive $\tau$ and $C$, and algebraically mixing if $\phi_{d}=Cd^{-r}$, for some positive $r$ and $C$.
Instantiating the bound of Theorem 3.9 to these two cases yields the following two corollaries.
Corollary 4.3.
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})}{n}+\frac{C}{n}+\sqrt{\frac{2\left(\tau\log n+1\right)\log\left(\frac{1}{\delta}\right)}{n}}\,.$
Up to a term linear in $\tau$ and some logarithmic factors, the above states that under the geometric mixing the same rates are achievable as in the i.i.d. setting. Roughly speaking, this amounts to saying that the effective sample size is a factor $\tau$ smaller than the original number of samples $n$, as long as generalization is concerned.
Corollary 4.4.
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})}{n}+C\left(1+\sqrt{\log(1/\delta)}\right)n^{-\frac{2r}{2(1+2r)}}\,.$
@@ 4.3 Multiplicative weights with delay
We now turn our attention to picking online strategies for the purpose of bounding the main term in the decomposition of the generalization error. We start by focusing on the classic exponential weighted average (EWA) algorithm (Vovk, 1990; Littlestone and Warmuth, 1994; Freund and Schapire, 1997). We fix a data-free prior $P_{1}\in\Delta_{\mathcal{W}}$ and a learning rate parameter $\eta>0$. We consider the updates
$P_{t+1}=\operatorname*{arg\ min}_{P\in\Delta_{\mathcal{W}}}\left\{\langle P,c_{t}\rangle+\frac{1}{\eta}\mathcal{D}_{KL}(P||P_{t})\right\},$
Combining the standard regret bound of EWA (see Appendix B.1) with Lemma 4.1 and Corollary 4.3 yields the result that follows.
Corollary 4.5.
@@
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathcal{D}_{KL}(P_{W_{n}|S_{n}}||P_{1})(\tau\log n+1)}{\eta n}+\frac{\eta}{2}+\frac{C}{n}+\sqrt{\frac{2\left(\tau\log n+1\right)\log\left(\frac{1}{\delta}\right)}{n}}\,.$
This results suggests that when considering geometric mixing processes, by applying a union bound over a well-chosen range of $\eta$ we recover the PAC-Bayes bound of McAllester (1998) up to a $O(\sqrt{\tau\log n})$ factor that quantifies the price of dropping the i.i.d. assumption. A similar result can be derived from Corollary 4.4 for algebraically mixing processes, leading to a bound typically scaling as $n^{-2r/(2(1+2r))}$.
@@ 4.4 Follow the regularized leader with delay
Corollary 4.6.
Suppose that $(Z_{t})_{t\in\mathbb{N}^{*}}$ is a geometric mixing process with constants $\tau,C>0$. Suppose that $\ell(w,z)\in[0,1]$ for all $w,z$. Assume there exists $B>0$ such that for all $t,||c_{t}||_{*}\leq B.$ Then, for any $P_{1}\in\Delta_{\mathcal{W}}$, for any $\delta>0$ with probability at least $1-\delta$ on the draw of $S_{n}$, uniformly for all $\mathcal{A}$,
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\left(h(P_{W_{n}|S_{n}})-h(P_{1})\right)(\tau\log n+1)}{\eta n}+\frac{\eta B^{2}}{2\alpha}+\frac{C}{n}+\sqrt{\frac{2\left(\tau\log n+1\right)\log\left(\frac{1}{\delta}\right)}{n}}.$
This generalization bound is similar to the bound of Theorem 9 of Lugosi and Neu (2023) up to a $O(\sqrt{\tau\log n})$ factor, when applying a union-bound argument over an appropriate grid of learning-rates $\eta$. In particular, this result recovers PAC-Bayesian bounds like those of Corollary 4.5 when choosing $h=\mathcal{D}_{\mathrm{KL}}(\cdot\|P_{1})$. We refer to Section 3.2 in Lugosi and Neu (2023) for more discussion on such bounds. As before, a similar result can be stated for algebraically mixing processes, with the main terms scaling as $n^{-2r/2(1+2r)}$ instead of $n^{-1/2}$.
@@ 5 Generalization bounds for dynamic hypotheses
Finally, inspired by the works of Eringis et al. (2022, 2024), we extend our framework to accommodate loss functions $\ell$ that rely not only on the last data point $Z_{t}$, but on the entire data sequence $\overline{Z}_{t}=\left(Z_{t},Z_{t-1},\dots,Z_{1}\right)$. Formally, we will consider loss functions of the form $\ell:\mathcal{W}\times\mathcal{Z}^{*}\rightarrow\mathbb{R}_{+}$ [^2] Here, $\mathcal{Z}^{*}$ denotes the disjoint union $\mathcal{Z}^{*}=\sqcup_{t\in\mathbb{N}}\mathcal{Z}^{t}$. and write $\ell(w,\overline{z}_{t})$ to denote the loss associated with hypothesis $w\in\mathcal{W}$ on sequence $\overline{z}_{t}\in\mathcal{Z}^{t}$. This consideration extends the learning problem to class of dynamical predictors such as Kalman filters, autoregressive models, or recurrent neural networks (RNNs), broadly used in time-series forecasting (Ariyo et al., 2014; Takeda et al., 2016). Specifically, if we think of $z_{t}=(x_{t},y_{t})$ as a data-pair of context and observation, in time-series prediction we usually not only rely on the context $x_{t}$ but also on the past sequence of contexts and observations $(x_{t-1},y_{t-1},\dots,x_{1},y_{1})$. As an example, consider $\ell(w,z_{t},\dots,z_{1})=\frac{1}{2}(y_{t}-h_{w}(x_{t},z_{t-1},\dots,z_{1}))^{2}$ where $h\in\mathcal{H}$ is a function class parameterized by $\mathcal{W}$. For this type of loss function a natural definition of the test error is:
@@
This is a generalization of Assumption 1 in the sense that taking $\ell(w,Z_{t},\dots,Z_{1})=\ell(w,Z_{t})$ simply amounts to requiring the same mixing condition as before. For our online-to-PAC conversion we consider the same framework as in Definition 3.4, except that now the cost function is defined as
@@
It easy to check that result of Lemma 3.10 still holds for this specific cost, and we can thus extend all the results of Section 4. We state the following adaptation of Theorem 3.9 below.
Theorem 5.1.
Assume $(Z_{t})_{t\in\mathbb{Z}}$ which satisfies Assumption 2 and consider a $d$-delayed online learning algorithm with regret bounded by $\mathrm{Regret}_{d,n}(P^{*})$ against any comparator $P^{*}$. Then, for any $\delta>0$, it holds with probability $1-\delta$:
$\mathrm{Gen}(\mathcal{A},S_{n})\leq\frac{\mathrm{Regret}_{d,n}(P_{W_{n}|S_{n}})}{n}+\phi_{d}+\sqrt{\frac{2d\log\left(\frac{1}{\delta}\right)}{n}}.$
@@
Assumption 3
There exists a decreasing non-negative sequence $(B_{d})_{d\in\mathbb{N}^{*}}$ such that, for any two sequences $\overline{z}_{t}=(z_{t},\dots,z_{i})$ and $\overline{z}_{t}^{\prime}=(z_{t}^{\prime},\dots,z_{j}^{\prime})$ of possibly different lengths that satisfy $z_{k}=z_{k}^{\prime}$ for all $k\in{t,\dots,t-d+1}$, we have $\left|\ell(w,\overline{z}_{t})-\ell(w,\overline{z}^{\prime}_{t})\right|\leq B_{d},$ for all $w\in\mathcal{W}$.
This condition can be verified for stable dynamical systems like autoregressive models, certain classes of RNNs, or sequential predictors that have bounded memory by design (see Eringis et al., 2022, 2024). The next assumption is a refinement of Assumption 1, adapted to the case where the loss function acts on blocks of $d$ data points $\overline{z}_{t-d+1:t}=(z_{t},z_{t-1},\dots,z_{t-d+1})$.
Assumption 4
Let $\overline{Z}_{t}=(Z_{t},\dots,Z_{1})$ be a sequence of data points and $\overline{Z}_{t}^{\prime}=(Z_{t}^{\prime},\dots,Z_{0}^{\prime},\dots)$ an independent copy of the same process. Then, there exists a decreasing sequence $(\beta_{d})_{d\in\mathbb{N}^{*}}$ non-negative real numbers such that the following is satisfied for all hypotheses $w\in\mathcal{W}$ and all $d\in\mathbb{N}^{*}$:
@@
This assumption can be verified whenever the loss function is bounded and the joint distribution of the data block $\overline{Z}_{t-d+1:t}$ satisfies a $\beta$-mixing assumption. This latter condition amounts to requiring that the conditional distribution of each data block given a block that trails $d$ steps behind is close to the marginal distribution in total variation distance, up to an additive term of $\beta_{d}$. The next proposition shows that these two simple conditions together imply that Assumption 2 holds, and the bound of Theorem 5.1 can be meaningfully instantiated for bounded-memory hypothesis classes deployed on mixing processes.
Proposition 5.2.